OpenStudy (anonymous):

Two systems of equations are shown below: System A System B 2x + y = 5 -10x + 19y = -1 -4x + 6y = -2 -4x + 6y = -2 Which of the following statements is correct about the two systems of equations?

OpenStudy (anonymous):

They will have the same solutions because the first equation of System B is obtained by adding the first equation of System A to 2 times the second equation of System A. They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 3 times the second equation of System A. The value of x for System B will be –5 times the value of x for System A because the coefficient of x in the first equation of System B is –5 times the coefficient of x in the first equation of System A. The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding –12 to the first equation of System A and the second equations are identical.

OpenStudy (anonymous):

System A 2x + y = 5 -4x + 6y = -2 System B -10x + 19y = -1 -4x + 6y = -2

OpenStudy (anonymous):

@tanmaykhandait